OpenStudy (anonymous):

how do i change this so i can graph it. 2x + 3y = -6

OpenStudy (whpalmer4):

No changing needed. Plug in \(x=0\) and find the value of \(y\). Make a dot at \((0,y)\). Plug in \(y=0\) and find the value of \(x\). Make a dot at \((x,0)\). Now draw a straight line through those points. You're done! However, if you wanted to do it the more tedious way, solve the equation for \(y\) in terms of \(x\): \[2x+3y=-6\]subtract \(2x\) from each side: \[2x-2x+3y = -6-2x\]\[3y=-6-2x\]divide each side by \(3\):\[\frac{3y}y = \frac{-6-2x}{3}\]\[y = -\frac{2}{3}x-2\]